A rectangles's length is 8 meters more than three times the width. The perimeter is 80 meters. What is the lenght?
8m
24m
32m
40m
I believe the answer is 24 or 32 but its tricky to me because of the 8 meters more part of the question. Also What is the difference between width and length. I thought they were the same. I understand that 2(w + h) = Perimeter
Thats why I think it is 32 or 24
Im sway to 24 when the 8 more is applied because it would mean that the height would have to be more than 8. HELP

Answer by scott8148(6628)

"A rectangles's length is 8 meters more than three times the width" ___ L = 3W + 8

"The perimeter is 80 meters" ___ 2L + 2W = 80 ___ L + W = 40 ___ W = 40 - L

substituting ___ L = 3(40 - L) + 8 ___ L = 120 - 3L + 8 ___ 4L = 128